A man is 4times as old as his son and five years ago the product of their ages is 234 find their present ages?
Answer by ankor@dixie-net.com(22740)

let m = the man's age
let s = the son's age
:
A man is 4 times as old as his son
m = 4s
:
five years ago the product of their ages is 234
(m-5)(s-5) = 234
replace m with 4s
(4s-5)(s-5) = 234
FOIL
4s^2 - 20s - 5s + 25 = 234
Combine to form a quadratic equation
4s^2 - 25s + 25 - 234 = 0
4s^2 - 25s - 209 = 0
Use the quadratic formula; a=4; b=-25; c=-209
I got a positive solution of:
s = 11 yr is the son's age
then
4(11) = 44 yrs is the man's age
:
:
see if the product (5 yrs ago) checks out: (11-5) * (44-5) = 234